A body executing shm has a max acceleration equal to 24m/sec^2 and max velocity equal to 16m/sec the amplitude of the shm is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Pallavi

let the equation of SHM is

x=Xsin(wt)

v = dx/dt=Xwcos(wt)

a=d2x/dt2 =-Xw2sin(wt)


given V max=Xw =16  ...........1

and a max =Xw2 =24  ........2

square equation and dived by 2

   X= 16*16/24

      =32/3


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed

solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin